A regular 30-sided polygon is inscribed in a circle with a radius of 8. Use the law of cosines to find an exact expression for the length of one of the sides.


Answer by jim_thompson5910(35256)

c^2 = a^2 + b^2 - 2ab*cos(C)

c^2 = 8^2 + 8^2 - 2*8*8*cos(12)

c^2 = 64 + 64 - 128*cos(12)

c^2 = 128 - 128*cos(12)

c = sqrt( 128 - 128*cos(12) )